AUG026
Hobart 26m
VSN: HOB+0076 start pos:
- Manual drive reset needed at beginning of experiment, drives operational at 1947UT (JS)
- First data at 1952UT missed scans 195-1800 – 195-1947 (JS)
- Windstowed between ~0200-0240
- onsource status slewing since UT13:19 (Liz)
Katherine
VSN: HOB+0093 start pos: 874.9 GB
- Experiment started OK (JS)
- eremote control not working when I arrived, apparently have just been using field system VNC since before 5pm local (UT08:00) (Liz)
Yaragadee
VSN: HOB+0041, start pos: 0 GB
- Experiment started OK (JS)
- Data recorded becoming consistently less than expected. Approx 200Gbyte at end. (liz)
- System monitor kept freezing-to fix, had to keep reopening (Liz)
~~DISCUSSION~~